# Quantum backreaction in an analogue black hole

### v1

@@ Abstract
We extend the Gross-Pitaevskii equation to incorporate the effect of quantum fluctuations onto the flow of a weakly interacting Bose-Einstein condensate. Applying this framework to an analogue black hole in a quasi-one-dimensional, transonic flow, we investigate how acoustic Hawking radiation back-reacts on the background condensate. Our results point to the emergence of stationary density and velocity undulations in the supersonic region (analogous to the black hole interior) and enable to evaluate the change in upstream and downstream Mach numbers caused by Hawking radiation. These findings provide new insight into the interplay between quantum fluctuations and analogue gravity in Bose-Einstein condensates.
@@
As first noted in Ref. [1], a one-dimensional weakly interacting Bose gas provides an excellent platform for realizing an analogy originally proposed by Unruh. According to this analogy, the interface between the supersonic and the subsonic regions of a transonic flow acts as a “sonic horizon” mimicking the event horizon of a gravitational black hole [2]. Within the Bose-Einstein condensate (BEC) framework, this line of research culminated in the experimental observation of analog Hawking radiation [3], manifested as nonlocal correlations in density fluctuations emitted from the acoustic horizon [4].
@@
From an experimental standpoint, the appeal of BEC platforms lies in their ultralow temperatures, their paradigmatic quantum nature, and the high degree of experimental control achievable in such systems. On the theoretical side, the Unruh analogy arises naturally: following Bogoliubov’s approach [5], it is customary to separate a classical field –representing the background flow that serves as an effective metric– from quantum fluctuations. However, the Bogoliubov approach is approximate, and it is natural to seek improvements by incorporating higher-order interactions among the classical and/or quantum components of the total field. This topic has been extensively studied in BEC physics, beginning with the pioneering work of Beliaev [6, 7].
@@
In the context of analogue gravity, this line of research is referred to as quantum backreaction, drawing the analogy with the study of how Hawking radiation affects the black hole metric [8, 9, 10, 11, 12]. In General Relativity, this problem is notoriously difficult and fundamentally constrained by the absence of a quantum theory of gravitation (for a recent review on the backreaction problem, see Ref. [13] and references therein). In contrast, the prospects are far more promising in BEC physics, where one effectively has a “theory of everything” since a single quantum field simultaneously governs the dynamics of both the effective metric and its quantum fluctuations. As a result, a number of theoretical studies addressed this issue (actually in a more general framework than the mere analogous Hawking radiation), both numerically [14, 15, 16, 17] and using perturbative methods [18, 19, 20, 21, 22, 23]. The latter closely parallel the semiclassical approach used in General Relativity: the equation governing the classical field (analogous to Einstein’s equations) is modified by a contribution from the expectation value of the quantum fluctuations, encompassed in the equivalent of a stress-energy tensor.
@@
However, in one-dimensional BECs, such approaches face technical challenges physically rooted in the Hohenberg-Mermin-Wagner theorem: at this dimensionality, the standard Gross-Pitaevskii description is particularly simple, but quantum phase fluctuations diverge at long wavelengths. A way around issues of this type has been proposed by Popov [24, 25, 26], who developed an amplitude-phase formalism to treat the long-wavelength degrees of freedom of a Bose field operator. This approach has since proven highly effective in BEC physics [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]; we will in the following combine it with a perturbative framework elaborated in Refs. [37, 38, 39, 40].
@@
The paper is organized as follows. In Sec. I, we introduce the model and the symmetry-breaking approach, which involves decomposing the quantum field operator into a classical order parameter and a (small) quantum correction. In Sec. II, we present an alternative, though not identical, separation that proves better suited to our objectives. Section III details our perturbative expansion. The leading order corresponds to the standard Gross-Pitaevskii equation (Sec. III.1). The next order – the Bogoliubov level – is presented in Sec. III.2, and is used in Sec. III.3 to compute the source terms that appear in the final step of our expansion. This last step, presented in Sec. III.4, leads to a modified equation for the classical order parameter that incorporates the effects of quantum fluctuations. The equations obtained in Sec. III.4 are general and apply to time- and space-dependent configurations in arbitrary dimensions. In Sec. IV we focus on stationary states. We first examine a uniform system in Sec. IV.1 to benchmark our approach, and then turn in Sec. IV.2 to the specific case of a one-dimensional analogue black hole—the primary system for which our theoretical framework has been devised. Our conclusions and prospects for future work are presented in Sec. V. Relevant results from earlier studies, as well as technical aspects, are summarized in the appendices. Appendices A and B present the acoustic black-hole configurations we consider and the associated quantum (outgoing or ingoing) modes. These modes are used in Appendix C as a basis for expanding the quantum fluctuation field, which makes it possible to explicitly compute source terms relevant to our backreaction equations. Appendix D details some steps of the computations presented in the main text. Our method is applied in Sec. IV.2 for a specific type of back hole configuration (the so-called “waterfall”), and Appendix E complements this discussion by presenting results for different analogue settings.
@@ I Self-consistent approach
The Bogoliubov approach amounts to separate in a Bose-condensed system a classical and a quantum contribution. This can be achieved by writing [41, 42]
@@
are known as the normal and anomalous averages, respectively. From these expressions, the average of Eq. (1) reads [39]
@@
The difference between Eqs. (8) and (1) reads [39]
@@ II Amplitude-phase formalism
where $\Theta(\boldsymbol{x},t)$ and $\rho(\boldsymbol{x},t)$ are the classical phase and density fields, respectively. The Hermitian field operators $\hat{\theta}(\boldsymbol{x},t)$ and $\hat{\eta}(\boldsymbol{x},t)$ describe quantum fluctuations of the phase and of the relative density, respectively, with $\langle\hat{\theta}\rangle=0=\langle\hat{\eta}\rangle$. The amplitude-phase decomposition can be compared with (2) by performing a series expansion of expression (10):
@@
In all the following it will suffice to keep in this expansion only the terms which have been explicitly written down in (11). This amounts to write
@@
$\Phi=\varphi(1+\langle\hat{A}\rangle),\quad\tilde{n}=|\varphi|^{2}\langle\hat{a}^{\dagger}\hat{a}\rangle,\quad\tilde{m}=\varphi^{2}\langle\hat{a}\hat{a}\rangle$, (15)
@@
The first of Eqs. (15) shows that the classical fields $\Phi(\boldsymbol{x},t)$ and $\varphi(\boldsymbol{x},t)$ are not identical, indicating that the separation between quantum and classical contributions in Eq. (2) does not exactly match the one in Eq. (10). The interest in the amplitude-phase formalism lies in the fact that, whereas in low dimension $\Phi$ is an ill defined quantity which obeys a singular equation, we will see in Sec. III.4 that the quantity $\varphi$ is well behaved. Accordingly, Eq. (8) which governs the dynamics of the classical field is written as
@@ III.1 Classical fields and Gross-Pitaevskii equation
Comparing Eqs. (17) and (18) makes it clear that the first effects of quantum fluctuations on $\delta\varphi(\boldsymbol{x},t)$ and $\delta\Phi(\boldsymbol{x},t)$ are of second order in $\hat{a}$ and $\hat{a}^{\dagger}$, which legitimates our discarding of contributions involving terms of third order in these fields in Sec. II, since they arise at the next order in perturbation. We may thus replace the first of Eqs. (15) by
@@ III.3 Averages of quantum fields
where the operators between colons are normal ordered and $\hat{\rho}$ is the density operator which exact expression (valid at all orders) is
@@
where Eq. (40) has been employed to obtain the last expression. The term between brackets in the third line of (46) is identical to the first average in the right hand side of the first line. It is known as the depletion of the condensate. It is ultraviolet convergent in dimensions 1, 2 and 3, but phase fluctuations make it infrared diverging in one dimension at zero temperature. It is this effect which forbids bona fide Bose-Einstein condensation in 1D. However, it has been shown in [43] that this divergence is position-independent and thus killed by the spatial derivative acting on $\langle\hat{A}\rangle$ in Eq. (53) below.
@@ III.4 Second order: backreaction for the classical fields
In the above expression the second line has been simplified –resulting in the final line– by observing that $\langle\hat{A}\rangle+\langle\hat{A}^{\dagger}\rangle=-\langle\hat{a}^{\dagger}\hat{a}\rangle$, as can be directly checked from expressions (14). This is an important step because in 1D the averages $\langle\hat{a}^{\dagger}\hat{a}\rangle$ and $\langle\hat{A}\rangle$ which appear in the second line of Eq. (48) are ill defined due to infrared divergent phase fluctuations, whereas the term $\langle\hat{a}^{\dagger}\hat{a}+\hat{a}\hat{a}\rangle$ is infrared regular. In 2D and 3D this source term is correctly taken into account by a proper renormalisation of the coupling constant which cancels an ultraviolet divergence in the anomalous average $\langle\hat{a}\hat{a}\rangle$, see Sec. IV below or also, e.g., Refs. [38, 30]. Hence, whereas Eq. (21) contains troublesome divergent terms, the only possible divergence in (48) comes from the $\langle\hat{A}\rangle$ term in the left hand side. However, it is acted upon by space and time derivatives which kill its diverging part as shown in [43] and verified below.
@@
$\begin{split}&\hbar(\partial_{t}+\boldsymbol{V}_{\!\!\scriptscriptstyle GP}\cdot\boldsymbol{\nabla})\delta\Theta-\frac{\hbar^{2}}{4m\rho_{\rm\scriptscriptstyle GP}}\boldsymbol{\nabla}\cdot\left(\rho_{\rm\scriptscriptstyle GP}\boldsymbol{\nabla}\frac{\delta\rho}{\rho_{\rm\scriptscriptstyle GP}}\right)\\ &+g\delta\rho=-\tfrac{1}{2}g\rho_{\rm\scriptscriptstyle GP}g^{(2)}-\frac{\hbar}{2}(\partial_{t}+\boldsymbol{V}_{\!\!\scriptscriptstyle GP}\cdot\boldsymbol{\nabla})\Re\langle\hat{\eta}\hat{\theta}\rangle\\ &-\frac{\hbar^{2}}{4m\rho_{\rm\scriptscriptstyle GP}}\boldsymbol{\nabla}\cdot\left[\rho_{\rm\scriptscriptstyle GP}\boldsymbol{\nabla}\left(\langle\hat{\theta}^{2}\rangle-\tfrac{1}{4}\frac{\delta(\boldsymbol{0}\,)}{\rho_{\rm\scriptscriptstyle GP}}+\tfrac{1}{4}g^{(2)}\right)\right].\end{split}$ (53)
At this point it is important to stress that the infrared divergence of the term $\langle\hat{\theta}^{2}\rangle$ in one dimension is killed by the spatial derivative in (53), a result which has been proven in Ref. [43]. Thus Eq. (52) can a priori be solved in any dimension, as illustrated below.
@@
Equations (53) and (54) are the main result of the present paper. They describe the dynamics of the leading order modification of the classical background flow induced by quantum fluctuations. They are valid even in a time-dependent flow for which the zeroth order Gross-Pitaevskii equation (18) and the first order Bogoliubov equations (23) (together with their solutions) are explicitly position and time-dependent. They constitute an extension of the results obtained by Mora and Castin in Ref. [30] to a possibly non-stationary flow in the presence of a background current. The amplitude-phase formalism makes it possible, without using a number conserving approach, to circumvent the infrared divergence problem encountered in other studies [22, 23].
@@ IV.1 Uniform and stationary configurations
This is known as the Hartree-Fock-Bogoliubov result for the chemical potential [37]. This expression differs from the Hugenholtz-Pines result [44] indicating that our approach, if pursued at higher order, would yield a gaped spectrum: in terms of the Hohenberg-Martin classification the present approach is conserving and not gapless, see discussions in [41, 37]. However, it is perfectly legitimate for our purpose, as illustrated in sections IV.1.1 and IV.1.2 (see also Refs. [38, 30]). Moreover, the violation of the Hugenholtz-Pines theorem is not necessarily prohibitive. A possible approach to recover a gapless spectrum—along with a corrected phonon velocity and damping rate—is to analyze the linear response of the classical field within a perturbative framework, as demonstrated by Giorgini in Ref. [40].

### v2

# Quantum backreaction in an analog black hole

@@ Abstract
We extend the Gross-Pitaevskii equation to incorporate the effect of quantum fluctuations onto the flow of a weakly interacting Bose-Einstein condensate. Applying this framework to an analog black hole in a quasi-one-dimensional, transonic flow, we investigate how acoustic Hawking radiation back-reacts on the background condensate. Our results point to the emergence of stationary density and velocity undulations in the supersonic region (analogous to the black hole interior) and enable to evaluate the change in upstream and downstream Mach numbers caused by Hawking radiation. These findings provide new insight into the interplay between quantum fluctuations and analog gravity in Bose-Einstein condensates.
@@
As first noted in Ref. [Garay2000], a one-dimensional weakly interacting Bose gas provides an excellent platform for realizing an analogy originally proposed by Unruh. According to this analogy, the interface between the supersonic and the subsonic regions of a transonic flow acts as a “sonic horizon” mimicking the event horizon of a gravitational black hole [Unruh1981]. Within the Bose-Einstein condensate (BEC) framework, this line of research culminated in the experimental observation of analog Hawking radiation [deNova2019], manifested as nonlocal correlations in density fluctuations emitted from the acoustic horizon [Balbinot2008].
@@
From an experimental standpoint, the appeal of BEC platforms lies in their ultralow temperatures, their paradigmatic quantum nature, and the high degree of experimental control achievable in such systems. On the theoretical side, the Unruh analogy arises naturally: following Bogoliubov’s approach [Bogoliubov1947], it is customary to separate a classical field –representing the background flow that serves as an effective metric– from quantum fluctuations. However, the Bogoliubov approach is approximate, and it is natural to seek improvements by incorporating higher-order interactions among the classical and/or quantum components of the total field. This topic has been extensively studied in BEC physics, beginning with the pioneering work of Beliaev [Beliaev1958a, Beliaev1958b].
@@
In the context of analog gravity, this line of research is referred to as quantum backreaction, drawing the analogy with the study of how Hawking radiation affects the black hole metric [Bardeen1981, York1984, Balbinot1986, Fabbri_2005, Hu2007]. In General Relativity, this problem is notoriously difficult and fundamentally constrained by the absence of a quantum theory of gravitation (for a recent review on the backreaction problem, see Ref. [delRio2025] and references therein). In contrast, the prospects are far more promising in BEC physics, where one effectively has a “theory of everything” since a single quantum field simultaneously governs the dynamics of both the effective metric and its quantum fluctuations. As a result, a number of theoretical studies addressed this issue (actually in a more general framework than the mere analogous Hawking radiation), both numerically [Robertson2018, Chatrchyan2018, Butera2019, Butera2023] and using perturbative methods [Balbinot2005, Balbinot2005a, Schultzhold2005, Liberati2020, Baak2022, Pal2024]. The latter closely parallel the semiclassical approach used in General Relativity: the equation governing the classical field (analogous to Einstein’s equations) is modified by a contribution from the expectation value of the quantum fluctuations, encompassed in the equivalent of a stress-energy tensor.
@@
However, in one-dimensional BECs, such approaches face technical challenges physically rooted in the Hohenberg-Mermin-Wagner theorem: at this dimensionality, the standard Gross-Pitaevskii description is particularly simple, but quantum phase fluctuations diverge at long wavelengths. A way around issues of this type has been proposed by Popov [Popov1971, Popov1972a, Popov1972b], who developed an amplitude-phase formalism to treat the long-wavelength degrees of freedom of a Bose field operator. This approach has since proven highly effective in BEC physics [Shevchenko1992, Petrov2000, Andersen2002, Mora2003, Petrov2004, Chiocchetta2013, Ji2015, Micheli2022, Duval2023, Micheli2024]. In the analysis below, we integrate it with the perturbative framework developed in Refs. [Griffin1996, Shi1998, Giorgini1998, Fedichev1998, Zaremba1999, Giorgini2000]. This set of studies incorporates corrections beyond the leading Gross–Pitaevskii order, capturing both the normal and anomalous averages [defined in Eq. (7)]. Such contributions are crucial for a consistent treatment of both quantum and finite-temperature fluctuations.
@@
The paper is organized as follows. In Sec. I, we introduce the model and the symmetry-breaking approach, which involves decomposing the quantum field operator into a classical order parameter and a (small) quantum correction. In Sec. II, we present an alternative, though not identical, separation that proves better suited to our objectives. Section III details our perturbative expansion. The leading order corresponds to the standard Gross-Pitaevskii equation (Sec. III.1). The next order – the Bogoliubov level – is presented in Sec. III.2, and is used in Sec. III.3 to compute the source terms that appear in the final step of our expansion. This last step, presented in Sec. III.4, leads to a modified equation for the classical order parameter that incorporates the effects of quantum fluctuations. The equations obtained in Sec. III.4 are general and apply to time- and space-dependent configurations in arbitrary dimensions. In Sec. IV we focus on stationary states. We first examine a uniform system in Sec. IV.1 to benchmark our approach, and then turn in Sec. IV.2 to the specific case of a one-dimensional analog black hole—the primary system for which our theoretical framework has been devised. Our conclusions and prospects for future work are presented in Sec. V. Relevant results from earlier studies, as well as technical aspects, are summarized in the appendixes. Appendixes A and B present the acoustic black-hole configurations we consider and the associated quantum (outgoing or ingoing) modes. These modes are used in Appendix C as a basis for expanding the quantum fluctuation field, which makes it possible to explicitly compute source terms relevant to our backreaction equations. Appendix D details some steps of the computations presented in the main text. Our method is applied in Sec. IV.2 for a specific type of back hole configuration (the so-called “waterfall”), and Appendix E complements this discussion by presenting results for different analogue settings.
@@ I Self-consistent approach
The Bogoliubov approach amounts to separate in a Bose-condensed system a classical and a quantum contribution. This can be achieved by writing [Hohenberg1965, Anderson1966]
@@
are known as the normal and anomalous averages, respectively. From these expressions, the average of Eq. (1) reads [Zaremba1999]
@@
The difference between Eqs. (8) and (1) reads [Zaremba1999]
@@ II Amplitude-phase formalism
where $\Theta(\boldsymbol{x},t)$ and $\rho(\boldsymbol{x},t)$ are the classical phase and density fields, respectively. The Hermitian field operators $\hat{\theta}(\boldsymbol{x},t)$ and $\hat{\eta}(\boldsymbol{x},t)$ describe quantum fluctuations of the phase and of the relative density, respectively, with $\langle\hat{\theta}\rangle=0=\langle\hat{\eta}\rangle$.
The perturbative expansions based on decomposition (2) and (10) are not equivalent. To make the difference explicit, we carry out a series expansion of expression (10):
@@
In all the following it will suffice to keep in this expansion only the terms which have been explicitly written down in (11). This amounts to including fluctuations up to quadratic order in $\hat{\eta}$ and $\hat{\theta}$ and to write
@@
[…]
$\tilde{n}=|\varphi|^{2}\langle\hat{a}^{\dagger}\hat{a}\rangle,\quad\tilde{m}=\varphi^{2}\langle\hat{a}\hat{a}\rangle$, (16)
@@
Eq. (15) shows that the classical fields $\Phi(\boldsymbol{x},t)$ and $\varphi(\boldsymbol{x},t)$ are not identical, indicating that the separation between quantum and classical contributions in Eq. (2) does not exactly match the one in Eq. (10). The interest in the amplitude-phase formalism lies in the fact that, whereas in low dimension $\Phi$ is an ill-defined quantity which obeys a singular equation, we will see in Sec. III.4 that the quantity $\varphi$ is well behaved. Accordingly, Eq. (8) which governs the dynamics of the classical field is written as
@@ III.1 Classical fields and Gross-Pitaevskii equation
Comparing Eqs. (18) and (19) makes it clear that the first effects of quantum fluctuations on $\delta\varphi(\boldsymbol{x},t)$ and $\delta\Phi(\boldsymbol{x},t)$ are of second order in $\hat{a}$ and $\hat{a}^{\dagger}$, which legitimates our discarding of contributions involving terms of third order in these fields in Sec. II, since they arise at the next order in perturbation. We may thus replace Eq. (15) by
@@ III.3 Averages of quantum fields
where the operators between colons are normal ordered and $\hat{\rho}$ is the density operator whose exact expression (valid at all orders) is
@@
where Eq. (41) has been employed to obtain the last expression. The term between brackets in the third line of (47) is identical to the first average in the right hand side of the first line. It is known as the depletion of the condensate. It is ultraviolet convergent in dimensions 1, 2 and 3, but phase fluctuations make it infrared diverging in one dimension at zero temperature. It is this effect which forbids bona fide Bose-Einstein condensation in 1D. However, it has been shown in [Anderson2015] that this divergence is position-independent and thus killed by the spatial derivative acting on $\langle\hat{A}\rangle$ in Eq. (54) below.
@@ III.4 Second order: backreaction for the classical fields
In the above expression the second line has been simplified –resulting in the final line– by observing that $\langle\hat{A}\rangle+\langle\hat{A}^{\dagger}\rangle=-\langle\hat{a}^{\dagger}\hat{a}\rangle$, as can be directly checked from expressions (14). This is an important step because in 1D the averages $\langle\hat{a}^{\dagger}\hat{a}\rangle$ and $\langle\hat{A}\rangle$ which appear in the second line of Eq. (49) are ill defined due to infrared divergent phase fluctuations, whereas the term $\langle\hat{a}^{\dagger}\hat{a}+\hat{a}\hat{a}\rangle$ is infrared regular. In 2D and 3D this source term is correctly taken into account by a proper renormalisation of the coupling constant which cancels an ultraviolet divergence in the anomalous average $\langle\hat{a}\hat{a}\rangle$, see Sec. IV below or also, e.g., Refs. [Giorgini1998, Mora2003]. Hence, whereas Eq. (22) contains troublesome divergent terms, the only possible divergence in (49) comes from the $\langle\hat{A}\rangle$ term in the left hand side. However, it is acted upon by space and time derivatives which kill its diverging part as shown in [Anderson2015] and verified below.
@@
$\begin{split}&\hbar(\partial_{t}+\boldsymbol{V}_{\!\!\scriptscriptstyle GP}\cdot\boldsymbol{\nabla})\left(\delta\Theta-\tfrac{1}{2}\Re\langle\hat{\eta}\hat{\theta}\rangle\right)+g(\delta\rho+\tfrac{1}{2}\rho_{\rm\scriptscriptstyle GP}g^{(2)})\\ &=\\ &\frac{-\hbar^{2}}{4m\rho_{\rm\scriptscriptstyle GP}}\boldsymbol{\nabla}\cdot\left[\rho_{\rm\scriptscriptstyle GP}\boldsymbol{\nabla}\left(\frac{\delta\rho}{\rho_{\rm\scriptscriptstyle GP}}-\langle\hat{\theta}^{2}\rangle+\frac{\delta(\boldsymbol{0})}{4\rho_{\rm\scriptscriptstyle GP}}-\frac{g^{(2)}}{4}\right)\right].\end{split}$ (54)
Equation (54) is an unsteady Bernoulli equation. Its classical version is usually obtained by performing a Madelung transform on the Gross-Pitaevskii equation. We consider here the next order which involves quadratic quantum contributions and the corresponding modifications of the density and phase of the classical field. In the first line it involves the classical modification of the phase (the velocity potential) together with a quantum fluctuation term ($\Re\langle\hat{\eta}\hat{\theta}\rangle$) and the interacting term resulting from the non-linear dependence of pressure on density, also affected by quantum fluctuations (the $g^{(2)}$ factor). The second line is the contribution of the so-called quantum pressure term which involves the modification $\delta\rho$ of the classical density and a quantum correction corresponding to the “quantum depletion” $\langle\hat{a}^{\dagger}\hat{a}\rangle=\langle\hat{\theta}^{2}\rangle-\tfrac{1}{4}\delta(\boldsymbol{0})/\rho_{\rm\scriptscriptstyle GP}+\tfrac{1}{4}g^{(2)}$, see Eq. (47). At this point it is important to stress that the infrared divergence of $\langle\hat{a}^{\dagger}\hat{a}\rangle$ in one dimension is killed by the spatial derivative in (54), a result which has been proven in Ref. [Anderson2015]. Thus Eq. (54) can a priori be solved in any dimension, as illustrated below.
@@
Equations (54) and (55) are the main result of the present paper. They describe the dynamics of the leading order modification of the classical background flow induced by quantum fluctuations. They are valid even in a time-dependent flow for which the zeroth order Gross-Pitaevskii equation (19) and the first order Bogoliubov equations (24) (together with their solutions) are explicitly position and time-dependent. They constitute an extension of the results obtained by Mora and Castin in Ref. [Mora2003] to a possibly non-stationary flow in the presence of a background current. The amplitude-phase formalism makes it possible, without using a number conserving approach, to circumvent the infrared divergence problem encountered in other studies [Baak2022, Pal2024]. Our treatment pertains to the so-called time-dependent Hartree-Fock-Bogoliubov scheme (see, e.g., discussions in Ref. [GNZ2009]). Note however, that we do not push the development as far as to include perturbatively coupled time-dependent equations for the condensate and noncondensate components of the Bose gas, as done by Giorgini in Ref. [Giorgini2000] for instance.
@@ IV.1 Uniform and stationary configurations
This is known as the Hartree-Fock-Bogoliubov result for the chemical potential [Griffin1996]. This expression differs from the Hugenholtz-Pines result [Hugenholtz1959] indicating that our approach, if pursued at higher order, would yield a gaped spectrum: in terms of the Hohenberg-Martin classification the present approach is conserving and not gapless, see discussions in [Hohenberg1965, Griffin1996]. However, it is perfectly legitimate for our purpose, as illustrated in sections IV.1.1 and IV.1.2 (see also Refs. [Giorgini1998, Mora2003]). Moreover, the violation of the Hugenholtz-Pines theorem is not necessarily prohibitive. A possible approach to recover a gapless spectrum—along with a corrected phonon velocity and damping rate—is to analyze the linear response of the classical field within a perturbative framework, as demonstrated by Giorgini in Ref. [Giorgini2000].